OpenStudy (anonymous):

vector a has a magnitude of 10 and a direction angle of 60 degrees. what is the magnitude and direction angle (as measured from the positive x-axis) of -A/2?

OpenStudy (aaronq):

i'm guessing that A is the original vector and you want to find \(\sf \dfrac{-A}{2}\)  so the magnitude is 10/2=5, the direction is 180+60=240\(^o\) from the positive x-axis in a counterclockwise direction.